OpenStudy (anonymous):

how would you integrate 5/3 t^(3/2)?

OpenStudy (nowhereman):

Using the power rule.

OpenStudy (anonymous):

As nowhereman has said,\[\int\limits_{}{}\frac{5}{3}t^{\frac{3}{2}}dt=\frac{5}{3}\frac{t^{\frac{3}{2}+1}}{\frac{3}{2}+1}+ c=\frac{5}{3}\frac{t^{5/2}}{\frac{5}{2}}+c=\frac{5}{3}.\frac{2}{5}t^{5/2}+c\]\[=\frac{2}{3}t^{5/2}+c\]

OpenStudy (anonymous):

Oh, thanks! I see this now! (just had to enlarge the screen!

OpenStudy (anonymous):

laurie if u are on today jun 28 before 12 midnite pls see my questions